OpenStudy (anonymous):

rewrite the given expression in terms of sum of difference of two angles: sin pi/7 cos pi/2 + cos pi/7 sin pi/2

zepdrix (zepdrix):

`Angle Sum Formula for Sine`:\[\large \sin \alpha \cos \beta+\cos \alpha\sin \beta \qquad=\qquad \sin\left(\alpha+\beta\right)\]